# Improving the efficiency of finite-time memory erasure with potential barrier shaping

### v4

@@ III.3 Analysis of the results with the detailed Jarzynski equality
We realize that in the classical set-up of the erasure mechanism in a symmetric bistable potential, the Landauer bound for the erasure of a classical bit of information is approached when the cycles are extremely slow, i.e., $\tau\to\infty$  [berut_2013, moupriya_2014, Moupriya_PhysRevE_2014]. However, it is understood that if the free energy change ($\Delta F$) for the process is estimated, the Landauer bound can be verified in terms of $\Delta F$ irrespective of the cycle duration $\tau$  [berut_2013, Buffoni_pre]. This is because, according to the thermodynamic principles, the free energy change corresponds to the lower limit of the work done for any process, in general. Now, the main question is how to get the difference in free energy between the initial and final states of the system for the erasure mechanism. To answer this question, we need to implement the Jarzynski equality, which allows us to calculate the free energy change between two states of a system in contact with a heat bath at temperature $T$. This measurement is done by employing the distribution of the work that is involved in the transformation between the states through nonequilibrium paths. The Jarzynski equality connects the change in free energy during any nonequilibrium process, in general, $\Delta F$, with the work done, $W$ [Jarzynski_prl_1997, Jarzynski_pre_1997, ChrisJarzynski_2004, sagwa_prl_jarjyanski_2010, buffoni2022spontaneous, pre_Williamson_2024], when both the initial and the final states are at equilibrium. The Jarzynski relation is expressed as follows,
@@
where $\beta=1/k_{B}T$. Now, we proceed to understand the free energy change for the erasure process through the application of the Jarzynski equality. For a complete erasure cycle, $f(t=0)=0$ and $f(t_{0}+t_{f})=f(t_{0})=0$. As this is a cyclic process, there will be no free energy change between the initial and final equilibrium states of the system, $F$ being a state function. i.e., $\Delta F=0$ or $\langle e^{-\beta W}\rangle=1$. This points towards the fact that the classical Jarzynski equality cannot be used to retrieve the Landauer bound in terms of the free energy change. This difficulty arises as the final state of the successful erasure cycle differs from the final equilibrium state. To estimate the free energy change from the work effects under these conditions, we need to consider the detailed Jarzynski equality [berut_2013, Kawai_2007, Vaikuntanathan_2009], which links the free energy change and nonequilibrium work when the final state is significantly different from the equilibrium state. We define an effective free energy change for the memory erasure process in which the memory ends in a unique desired state, as $\Delta F_{\text{effective}}$. This effective free energy is related to the nonequilibrium work through the relation [moupriya_2014, Moupriya_PhysRevE_2014]
@@
where $\langle\cdot\rangle_{\to 1}$ denotes the average taken over trajectories that end in the target state at the end of the erasure cycle. For our present setup, this desired final state is the right well or the memory state 11. We divide this into two sub-processes, as in  [berut_2013, moupriya_2014]: one process being the transitions of particles from the left to the right well, and the other considers the movements of the particles in the right well, which were already there at the start of the erasure protocol. In other words, the first sub-process can be described as the transition from the memory state 0 to $1$ during the erasure cycle, and the second one defines the scenario where the memory state remains unaltered as $1$ at the beginning and the end of the erasure protocol. The values of the average exponential work functions for these two sub-processes are defined as:
@@
where $\langle\cdot\rangle_{i\to j}$ is the average value of the function for trajectories of particles that initially start in the $i$th memory state and end in the $j$th memory state after the erasure protocol. With this definition, Eq. 11 can be rewritten as [moupriya_2014],
@@
The factor $1/2$ appears because of equal distribution in both wells after the thermalization. We have plotted $A_{01}$, $A_{11}$, and $\Delta F$ for various asymmetry parameters $(c)$ as functions of $\tau$. We found that for the symmetric case $(c=1)$, the sum of $A_{01}+A_{11}$ is close to $1$, which has already been reported [berut_2013, moupriya_2014], and this value corresponds to $\langle e^{-\beta W}\rangle_{\to 1}=1/2$, i.e., the free energy change is $\Delta F_{\text{effective}}=D\ln 2$. This establishes the fact that the Landauer bound can be recovered in terms of free energy change. The results have been shown in Fig. 14(a). However, we observed that in the asymmetric cases, the sum is always greater than $1$ for all degrees of asymmetry, because the system always tries to maximize the entropy and prefers the wider potential well. Therefore, the free energy calculation is
@@
where $p_{01}=\frac{N_{01}}{N_{01}+N_{11}}$ and $p_{11}=\frac{N_{11}}{N_{01}+N_{11}}$ are the relative fractions of each type of transition, conditioned on the final memory state being “1”. The free energy change falls below the Landauer bound, as shown in Figs. 14(b)-(e) for four different asymmetric cases studied.
@@
Figure 14: Plots of $A_{01}$, $A_{11}$, $A_{01}+A_{11}$ and $\Delta F$ versus $\tau$ for five different values of $c$ ($c=1$, $c=0.74$, $c=0.56$, $c=0.34$, and $c=0.20$), with fixed barrier height $h=1$, noise strength $D=0.15$, and barrier lowering parameters $\omega=2\pi/\tau$, and $Z=0.5$. The erasure cycles follow a ramp-up and ramp-down time ratio of $\tau_{1}:\tau_{2}=19:1$ for the tilting force. The corresponding amplitude of the tilting force are $A=5.0,0.9,0.75,0.65,$ and $0.55$, respectively. In all cases, the erasure protocols start with the initial equilibrium distribution of the memory states.
[…]
@@
There is one crucial aspect to mention at this point. In the case of the interpretation of the Landauer bound in terms of the effective free energy change for the erasure mechanism through the application of the detailed Jarzynski equality, we consider the equilibrium distribution of the initial memory states. This has been done for all systems under study, symmetric and asymmetric systems, with varied levels of asymmetries. This is because the implementation of the detailed Jarzynski equality will be valid even if the final state is a nonequilibrium state; however, the initial state has to be at equilibrium. The effective free energy change cannot be defined if both the initial and final states are not at equilibrium. Therefore, we depict the analyses of the effect of the asymmetry of the intrinsic potential on the effective free energy change for the erasure process, where we consider equilibrium initial conditions for all cases examined.
@@
Figure 15: Plot of $\Delta F$ versus $c$ at a fixed $\tau=1300$, with a constant barrier height $h=1$, noise strength $D=0.15$, and barrier lowering parameters $\omega=2\pi/\tau$, and $Z=0.5$. The erasure cycles follow a ramp-up and ramp-down time ratio of $\tau_{1}:\tau_{2}=19:1$ for the tilting force. In all cases, the erasure protocols start with the initial equilibrium distribution of the memory states.
@@ IV Conclusion and Outlook
In our current work, we focus on some crucial aspects linked to the fundamental process of information erasure where the information is stored as the classical bits of binary memory, 0 and $1$. The novel aim of the study is to explore the possibility of improving the performance of the erasure process, both in terms of the success rate and the amount of essential heat evolution and work-done associated with it. These are the critical factors that define the overall efficiency with which tasks are carried out in digital devices as the erasure steps are indispensable for their necessary operations. The goal is to increase the success rate and to decrease the extent of heat released to the environment, for the erasure mechanism, which would effectively improve the performance of the devices. The limit for the evolved heat for the erasure of a classical bit of memory is set by the Landauer bound of $k_{B}T\ln 2$, which is approached for an erasure protocol of infinite duration. A greater amount of heat is released for any finite-time erasure process. In the case of the small systems, the same description applies to the average of the quantity, heat. This is understood through the design of the erasure mechanism when the two memory states are modeled in a symmetric bistable potential, and they are separated by a symmetric barrier. We address the objectives of the present work, by carrying out a systematic analysis of modulating the degree of asymmetry of the underlying potential barrier separating the two binary memory states 0 and $1$. This asymmetry is implemented through different widths of the two potential wells of the bistable system and the asymmetric shape of the potential barrier partitioning them, although the two minima of the potential wells representing the memory states are energetically equivalent. We aim to scrutinize whether the asymmetry in the potential wells and potential barrier has any significant impact on the improvements of the effectiveness of the erasure protocol.
@@
It has been found recently [Chupeau_2020, innerbichler_2020] that the rate of the diffusive transport can be increased by optimized tuning of the potential barrier shape, which modulates the potential to different degrees of asymmetry. The key observation is that the speed of the barrier crossing process enhances for the asymmetric barrier shapes as compared to the symmetric case when all other conditions remain the same. Also, a larger extent of asymmetry has greater significance on the increase in the rate. We proceed to investigate whether this fundamental finding on the transition processes influences the erasure phenomenon that can be viewed as a driven barrier-crossing mechanism. Our study reveals that the erasure process becomes more favorable in terms of the success rate when the potential is more asymmetric in nature. To explain, there is a minimum amplitude of the erasure force, i.e., of the external linear drive which allows for a desired threshold of the success rate ($\sim 99\%$) for the erasure process. This optimum value of the forcing amplitude has a lesser magnitude when the degree of asymmetry in the potential is higher. This implies that the target level of the success rate or the erasure rate can be achieved by the application of a lower amount of the external force making the process feasible under moderate input conditions for asymmetric cases. This result can be considered as one of the significant outcomes of the present analysis as it indicates a direct and methodical way of increasing the success rate of the process by raising the measure of asymmetry in the potential. The increased success rate interpreted as the facilitated transition over the barrier, which essentially forms the erasure mechanism can be regarded as the improved kinetics of the process in asymmetric potential setups. To depict the effect of this asymmetry on the energetic aspects, we concentrate on the relevant thermodynamic quantities, such as the average work done, evolved heat and free energy change related to the erasure process. As the erasure protocols are set up as cyclic processes, i.e., the initial and final values of the external forces possess the same value, the numerical value of the average work done becomes equivalent to the average heat evolution for erasure following the thermodynamic principles. Our study leads to a very crucial understanding that the average work done or the released heat can reach below the Landauer bound for the finite-time erasure processes under pertinent conditions in the case of the asymmetric setups as described. This has also been justified by the calculated free energy change, which defines the lower limit of the average work done, through the application of the detailed Jarzynski equality. The deviation from the Landauer bound in the case of the asymmetric memory has been reported before [Konopik_2020_epl, Gavrilov_2016, talukdar2018designing]. Our work supports the previous findings. We also compare the crucial aspects of the investigations to analyze the effect of the equilibrium and nonequilibrium distributions of the initial memory states on them, wherever applicable.
@@
Here, we mention that the originality of our present study lies in the fact that we provide a thorough quantitative interpretation of the deviation from the Landauer bound depending on the degree of asymmetry in the underlying potential. We emphasize that it is possible to propose appropriate designs of the systems that can improve the operations of the erasure mechanisms in terms of the thermodynamic costs when the initial memory states have an asymmetric equilibrium or nonequilibrium origin. However, we retrieve the Landauer bound under the conventional setup with initially symmetric and equal distribution of memory states. This comparative study gives us the scope of proper quantitative understanding of improving the suggested erasure protocols from the perspective of energetics. Also, our analyses unveil another very vital fact that the success of the erasure process can be enhanced through the tuning of the asymmetry factor governing the dynamics. Therefore, it can be considered that the current work defines a systematic path and develops a detailed comprehension for the improvement of the efficiency of the fundamental process of erasure. We believe that the knowledge acquired from the present study would guide the design of more efficient erasure mechanisms. This will improve the performance of the digital devices and consequently have symbolic and beneficial effects on developing technologies.

### v5

@@ III.3 Analysis of the results with the detailed Jarzynski equality
We realize that in the classical set-up of the erasure mechanism in a symmetric bistable potential, the Landauer bound for the erasure of a classical bit of information is approached when the cycles are extremely slow, i.e., $\tau\to\infty$  [73, 74, 75]. However, it is understood that if the free energy change ($\Delta F$) for the process is estimated, the Landauer bound can be verified in terms of $\Delta F$ irrespective of the cycle duration $\tau$  [73, 50]. This is because, according to the thermodynamic principles, the free energy change corresponds to the lower limit of the work done for any process, in general. Now, the main question is how to get the difference in free energy between the initial and final states of the system for the erasure mechanism. To answer this question, we need to implement the Jarzynski equality, which allows us to calculate the free energy change between two states of a system in contact with a heat bath at temperature $T$. This measurement is done by employing the distribution of the work that is involved in the transformation between the states through nonequilibrium paths. The Jarzynski equality connects the change in free energy during any nonequilibrium process, in general, $\Delta F$, with the work done, $W$ [76, 77, 78, 79, 80, 81], when both the initial and the final states are at equilibrium. The Jarzynski relation is expressed as follows,
@@
where $\beta=1/k_{B}T$. Now, we proceed to understand the free energy change for the erasure process through the application of the Jarzynski equality. For a complete erasure cycle, $f(t=0)=0$ and $f(t_{0}+t_{f})=f(t_{0})=0$. As this is a cyclic process, there will be no free energy change between the initial and final equilibrium states of the system, $F$ being a state function. i.e., $\Delta F=0$ or $\langle e^{-\beta W}\rangle=1$. This points towards the fact that the classical Jarzynski equality cannot be used to retrieve the Landauer bound in terms of the free energy change. This difficulty arises as the final state of the successful erasure cycle differs from the final equilibrium state. To estimate the free energy change from the work effects under these conditions, we need to consider the detailed Jarzynski equality [73, 82, 83], which links the free energy change and nonequilibrium work when the final state is significantly different from the equilibrium state. We define an effective free energy change for the memory erasure process in which the memory ends in a unique desired state, as $\Delta F_{\text{effective}}$. This effective free energy is related to the nonequilibrium work through the relation [74, 75]
@@
where $\langle\cdot\rangle_{\to 1}$ denotes the average taken over trajectories that end in the target state at the end of the erasure cycle. For our present setup, this desired final state is the right well or the memory state 11. We divide this into two sub-processes, as in  [73, 74]: one process being the transitions of particles from the left to the right well, and the other considers the movements of the particles in the right well, which were already there at the start of the erasure protocol. In other words, the first sub-process can be described as the transition from the memory state 0 to $1$ during the erasure cycle, and the second one defines the scenario where the memory state remains unaltered as $1$ at the beginning and at the end of the erasure protocol. The values of the average exponential work functions for these two sub-processes are defined as:
@@
where $\langle\cdot\rangle_{i\to j}$ is the average value of the function for trajectories of particles that initially start in the $i$th memory state and end in the $j$th memory state after the erasure protocol. With this definition, Eq. 11 can be rewritten as [74],
@@
The factor $1/2$ appears because of equal distribution in both wells after the thermalization. We have plotted $A_{01}$, $A_{11}$, and $\Delta F_{\textrm{effective}}$ for various asymmetry parameters $(c)$ as functions of $\tau$. We found that for the symmetric case $(c=1)$, the sum of $A_{01}+A_{11}$ is close to $1$, which has already been reported [73, 74], and this value corresponds to $\langle e^{-\beta W}\rangle_{\to 1}=1/2$, i.e., the free energy change is $\Delta F_{\text{effective}}=D\ln 2$. This establishes the fact that the Landauer bound can be recovered in terms of free energy change. The results have been shown in Fig. 15(a). However, we observed that in the asymmetric cases, the sum is always greater than $1$ for all degrees of asymmetry, because the system always tries to maximize the entropy and prefers the wider potential well. Therefore, the free energy calculation is
@@
where $p_{01}=\frac{N_{01}}{N_{01}+N_{11}}$ and $p_{11}=\frac{N_{11}}{N_{01}+N_{11}}$ are the relative fractions of each type of transition, conditioned on the final memory state being “1”. The free energy change falls below the Landauer bound, as shown in Figs. 15(b)-(e) for four different asymmetric cases studied. It is noticed that the value of the estimated $\Delta F_{\textrm{effective}}$ is specific for each value of $c$. It increases with increasing $c$ and ultimately converges to the Landauer bound for the symmetric potential with $c=1$. These findings have been demonstrated in Fig. 16 for a specific value of the erasure cycle duration $\tau$. Now, very importantly, we identify that the $\langle W\rangle$ or $\langle Q\rangle$ approaches the calculated value of the corresponding $\Delta F_{\textrm{effective}}$ for larger durations of the erasure cycles, irrespective of the value of the asymmetry parameter $c$ (Fig.11). It is also observed in this plot (Fig.11) that the values of the $\Delta F_{\textrm{effective}}$ are more or less constant, for all values of $\tau$, with insignificant fluctuations, as expected, for a given value of $c$. This allows us to establish a bound on the evolved heat or associated work for the bit erasure process for asymmetric systems as well. We also compare $\langle W\rangle$ or $\langle Q\rangle$ with $\Delta F_{\textrm{effective}}$ in the probability distribution plots of $W$ and $Q$ for some representative values of $c$ (Figs.10(a), (c), and (e)). The decreasing magnitudes of this limiting value, interpreted in terms of $\Delta F_{\textrm{effective}}$, for higher degrees of asymmetry in the potential, establish the fact that the erasure protocols designed with asymmetric potentials can achieve lower thermodynamic costs. Our systematic analyses develop a quantitative understanding of this advantage that can be regulated through the tuning of the degree of asymmetry of the underlying potential. We emphasize that the outcomes of this particular investigation regarding $\Delta F_{\textrm{effective}}$ appear to be very significant. We make this remark as this exploration allows us to suggest a limit for the average values of the work or heat involvement for the erasure processes in small systems when the Landauer bound is supposed not to hold due to the intrinsic asymmetry of the governing potential.
@@
Figure 15: Plots of $A_{01}$, $A_{11}$, $A_{01}+A_{11}$ and $\Delta F_{\textrm{effective}}$ versus $\tau$ for five different values of $c$ ($c=1$, $c=0.74$, $c=0.56$, $c=0.34$, and $c=0.20$), with fixed barrier height $h=1$, noise strength $D=0.15$, and barrier lowering parameters $\omega=\pi/\tau$, and $Z=0.5$. The erasure cycles follow a ramp-up and ramp-down time ratio of $\tau_{1}:\tau_{2}=19:1$ for the tilting force. The corresponding amplitude of the tilting force are $A=5.0,0.9,0.75,0.65,$ and $0.55$, respectively. In all cases, the erasure protocols start with the initial equilibrium distribution of the memory states.
@@
There is one crucial aspect to mention at this point. In the case of the interpretation of the lower limit of the work or heat involvement in terms of the effective free energy change for the erasure mechanism through the application of the detailed Jarzynski equality, we consider the equilibrium distribution of the initial memory states. This has been done for all systems under study, symmetric and asymmetric systems, with varied levels of asymmetries. This is because the implementation of the detailed Jarzynski equality will be valid even if the final state is a nonequilibrium state; however, the initial state has to be at equilibrium. The effective free energy change cannot be defined if both the initial and final states are not at equilibrium. Therefore, we depict the analyses of the effect of the asymmetry of the intrinsic potential on the effective free energy change for the erasure process, where we consider equilibrium initial conditions for all cases examined. As the calculation of the effective free energy bound is not possible for all cases (with initial nonequilibrium distributions of the memory states), we preserve the Landauer bound line as a reference for comparison with the $\langle W\rangle$ or $\langle Q\rangle$ in Figs.10, 11, and 12. Although the calculations of $\Delta F_{\textrm{effective}}$ are not possible for the cases with nonequilibrium distributions of the initial memory states, we can comment on considering this quantity as a general lower bound for $\langle W\rangle$ or $\langle Q\rangle$ for the erasure processes regardless of the initial conditions. This is because when we scrutinize the values of $\langle W\rangle$ or $\langle Q\rangle$ for two distinct initial distributions, keeping all other factors the same, we find that the values are slightly higher for the nonequilibrium starting condition cases. The corresponding values become equal only for the symmetric system, as in that case, the initial equilibrium distribution and a 50:50 distribution of the memory states are identical. Therefore, we validate that the values of $\langle W\rangle$ or $\langle Q\rangle$ do not reach below $\Delta F_{\textrm{effective}}$ in any cases for the erasure mechanisms under study. Consequently, we can say that the quantity sets a general lower limit for the work or heat of erasure of a classical bit of information when the initial memory states can have asymmetric as well as nonequilibrium characteristics.
III.4 Validity in the parameter regime
Following the thorough analyses of the erasure process, exploiting the asymmetry of the governing potential, through the success rate, work (heat), and free energy calculations, we now explicitly discuss the justification of the choice of parameters in these studies. This will give a definite idea about the range of parameter values, linked to the system and dynamics, for which the above interpretations will be valid.
First, we discuss the rationale regarding the selected parameter values in the stage of the preparation of the memory states before the start of the erasure protocol. Throughout the study, the time-independent barrier height of the potential has been kept fixed at $h=1$, irrespective of the value of the asymmetry parameter $c$. In most cases, the value of $D$ is considered to be $0.15$. Therefore, compared to the barrier height, the noise strength satisfies the condition of a weak noise regime for the systems under study. This property of the environmental noise is required to certify the stability of the designed memory states. As the noise strength is sufficiently low compared to the barrier height of the potential, the transitions between the two wells, i.e., between the two memory states, are rare. This aspect takes care of the fact that the two types of memory states are well-defined, separated by a significant magnitude of the energy barrier. In the case of the studies with the initial thermalized distributions, the systems are first allowed to get equilibrated with the heat bath for a certain range of time. This time span is adequate for the systems to reach their steady state distributions as no significant changes in the distributions are observed if the systems evolve for a longer time. This fact has been verified for the system with the symmetric potential as well as for all asymmetric systems with a range of $c$ values considered in our studies. This examination suggests the robustness of the initial thermalized distribution and, consequently, the preparation of the memory states at the start of the protocol for the concerned cases. For the analyses with initial nonequilibrium distributions, we start with the $50:50$ distribution of the memory states for the systems with all values of $c$. In this case, we do not let the system get thermalized with the heat bath. Therefore, the thermal noise does not affect the preparation of the memory states for these circumstances.
Next, we scrutinize the chosen parameter values during the operation of the erasure cycles, to remark on the reliability of the selection of their magnitudes in constructing apt erasure mechanisms. The parameter $Z$ present in the barrier-lowering force (Eq. 4) determines the minimum height of the potential barrier during the erasure mechanism. As its value has been considered to be equal to $0.5$ for all cases studied, the lowest value that the barrier height can attain is $0.5$. Therefore, we suggest that a reasonable value of the barrier height to noise ratio (greater than $3$, for our specific choice of parameters) is maintained even for the situation with the least barrier height. This implies that the effect of the noise will not be dominant in the overall dynamics, so that the roles of the underlying potential and the erasure forces remain obscured during the process. At the same time, we mention that the barrier-lowering protocol is implemented in the erasure mechanism to allow for a greater amount of thermally activated transfer between the two states as compared to the unperturbed system [16, 15]. In the latter case, the distinction between the two memory states is more rigid and requires a stricter restriction on the observed overall stability of the two states. With reference to these facts, we point out that the required conditions are maintained throughout our studies, both for the initial conditions and during the erasure protocol. This ultimately verifies that the resulting erasure phenomenon captures the integrated effect of the underlying potential, environmental noise and the erasure forces appropriately. We also discuss the choice of the value of $\omega$, which is related to the tilting force time period $\tau$ for finite-time erasure as $\pi/\omega$. This consideration guarantees that the barrier-lowering force arrives at its initial value 0 at the end of the cycle, following the appropriate protocol. In the majority of the studies, the value of $\omega$ has been taken to be equal to one, which corresponds to $\tau=314$. This value of $\tau$ is comparable to the thermal relaxation time-scale of the systems, which has been discussed in the context of the initial thermalization of the two memory states. We also performed studies for varying values of $\tau$ and the magnitude of $\omega$ has been defined accordingly.
Finally, we indicate that the observed variation of the success rate versus $D$ plot (Fig.7(a)) is very informative to comment on the validation of the considered value of $D$ in comparison to the barrier height $h$. We detect that the success rate remains above a high threshold value (approximately above $95\%$) for all values of $c$, i.e., for all asymmetric and symmetric systems. This indicates that the erasure process works efficiently within this range of the noise strength for the class of systems under study. As we perform most of our analyses, considering $D$ value equal to $0.15$, we suggest that the results obtained from the present study correspond to effective erasure mechanisms. Also, we imply that the overall interpretation of the present work regarding the improvements of the erasure process in terms of both kinetic and thermodynamic factors will be valid within this weak noise range.
@@ III.4 Validity in the parameter regime
Figure 16: Plot of $\Delta F_{\textrm{effective}}$ versus $c$ at a fixed $\tau=1300$, with a constant barrier height $h=1$, noise strength $D=0.15$, and barrier lowering parameters $\omega=\pi/\tau$, and $Z=0.5$. The erasure cycles follow a ramp-up and ramp-down time ratio of $\tau_{1}:\tau_{2}=19:1$ for the tilting force. In all cases, the erasure protocols start with the initial equilibrium distribution of the memory states.
@@ IV Conclusion and Outlook
In our current work, we focus on some crucial aspects linked to the fundamental process of information erasure where the information is stored as the classical bits of binary memory, 0 and $1$. The novel aim of the study is to explore the possibility of improving the performance of the erasure process, both in terms of the success rate and the amount of essential heat evolution and work-done associated with it. These are the critical factors that contribute to the overall efficiency with which tasks are carried out in digital devices as the erasure steps are indispensable for their necessary operations. The goal is to increase the success rate and to decrease the extent of heat released to the environment, for the erasure mechanism, which would effectively improve the performance of the devices. The limit for the evolved heat for the erasure of a classical bit of memory is set by the Landauer bound of $k_{B}T\ln 2$, which is approached for an erasure protocol of infinite duration. A greater amount of heat is released for any finite-time erasure process. In the case of the small systems, the same description applies to the average of the quantity, heat [15]. This is understood through the design of the erasure mechanism when the two memory states are modeled in a symmetric bistable potential, and they are separated by a symmetric barrier. We address the objectives of the present work, by carrying out a systematic analysis of modulating the degree of asymmetry of the underlying potential barrier separating the two binary memory states 0 and $1$. This asymmetry is implemented through different widths of the two potential wells of the bistable system and the asymmetric shape of the potential barrier partitioning them, although the two minima of the potential wells representing the memory states are energetically equivalent. For this purpose, we consider the potential as described in Eq. 2. Despite the distinct mathematical form of this bistable potential and the particular parametrization, the potential function for asymmetric cases can account for the generic asymmetry in the dynamics through the introduction of an entropic bias between the two states, whereas keeping their energies similar  [54]. Therefore, this choice of potential guarantees the energy equivalence of the two binary memory states. Also, at the same time, it allows us to investigate the effect of the asymmetry of the governing potential on the erasure process, in general. We systematically scrutinize this aspect of the asymmetry in the potential form to find out any significant impact on the improvements of the effectiveness of the erasure protocol.
@@
It has been found recently [84, 54] that the rate of the diffusive transport can be increased by optimized tuning of the potential barrier shape, which modulates the potential to different degrees of asymmetry. The key observation is that the speed of the barrier crossing process enhances for the asymmetric barrier shapes as compared to the symmetric case when all other conditions remain the same. Also, a larger extent of asymmetry has greater significance on the increase in the rate. We proceed to investigate whether this fundamental finding on the transition processes influences the erasure phenomenon that can be partly viewed as a driven barrier-crossing mechanism. This is because the proportion of the memories that were initially at a different state other than the target state, are required to be transferred to the desired state at the end of the erasure process. In our physical setup, this is realized as the movement of the Brownian particles to the right well (memory state $1$), if they were primarily present at the left well (memory state 0), to account for the successful erasure. Our study reveals that the erasure process becomes more favorable in terms of the success rate when the potential is more asymmetric in nature. To explain, there is a minimum amplitude of the erasure force, i.e., of the external linear drive which allows for a desired threshold of the success rate ($\sim 99\%$) for the erasure process. This optimum value of the forcing amplitude has a lesser magnitude when the degree of asymmetry in the potential is higher. This implies that the target level of the success rate can be achieved by the application of a lower amount of the external force making the process feasible under moderate input conditions for asymmetric cases. This result can be considered as one of the significant outcomes of the present analysis as it indicates a direct and methodical way of increasing the success rate of the process by raising the measure of asymmetry in the potential. The increased success rate resulting from the facilitated transition over the barrier in the appropriate cases, can be regarded as the improved kinetics of the process in asymmetric potential setups. To depict the effect of this asymmetry on the energetic aspects, we concentrate on the relevant thermodynamic quantities, such as the average work done, evolved heat and the effective free energy change related to the erasure process. We compare the crucial aspects of the investigations to analyze the effect of the equilibrium and nonequilibrium distributions of the initial memory states on them, wherever applicable. As the erasure protocols are set up as cyclic processes, i.e., the initial and final values of the external forces possess the same value, the numerical value of the average work done becomes equivalent to the average heat evolution for erasure following the thermodynamic principles. Our study leads to a very crucial understanding that the average work done or the released heat can reach below the Landauer bound for the finite-time erasure processes under pertinent conditions in the case of the asymmetric setups as described. This has also been justified by the calculated effective free energy change, which defines the lower limit of the average work done, through the application of the detailed Jarzynski equality. The estimation of this effective free energy change, with the aid of the detailed Jarzynski equality, leads us to define a general bound for the average work done or evolved heat for the erasure processes linked to asymmetric systems. This finding is important as we identify a quantity, $\Delta F_{\textrm{effective}}$, which provides an idea about the limiting value of $\langle W\rangle$ or $\langle Q\rangle$ when the Landauer bound does not hold due to the asymmetric nature of the memory states. Here, we point out that deviation from the Landauer bound in the case of the asymmetric memory has been reported before [48, 55, 85] and our work supports the previous findings. Moreover, we perform and present a detailed analysis through which the departure from the Landauer bound can be understood quantitatively with respect to the degree of asymmetry of the system. Our thorough study results in establishing the effective free energy change as the lower limit for the average work done and heat released for the erasure mechanism. This becomes one of the important measures which can be employed to interpret the comprehensive variations of the energetic aspects of erasure of information for asymmetric systems as compared to the idealized case with symmetric memories.
Now, we express our views on the applicability and interpretations of the above-mentioned results for the erasure process in a general asymmetric framework. As our choice of the potential can validate the results of earlier studies, both in the theoretical and experimental cases [52, 55], we suggest that the considered potential function can be regarded to represent a general setup for the described type of asymmetry. The advantage of the present mathematical form of the potential is that the degree of asymmetry can be directly tuned through the asymmetry parameter $c$. The work [84] suggests that it is possible to understand the actions of the specified type of asymmetric potential through the design of experiments involving intricate tools of microfluidics and optical tweezers. Therefore, there remains further scope for verifying our numerical results on the kinetic and thermodynamic aspects of erasure of information in asymmetric systems. Here, the mentioned asymmetry characterizes the different phase-space volumes of the two types of binary memory states involved in the erasure mechanism.
@@
Here, we refer to that the originality of our present study lies in the fact that we provide a thorough quantitative interpretation of the improved outputs for the erasure processes designed in appropriate asymmetric setups. This advancement is realized from both the perspective of kinetics; in terms of the success rate, and energy costs; estimated through $\langle W\rangle$ or $\langle Q\rangle$, and $\Delta F_{\textrm{effective}}$. Our thorough numerical study quantifies the deviation from the Landauer bound for $\langle W\rangle$ or $\langle Q\rangle$ depending on the degree of asymmetry in the underlying potential. The detailed analyses lead to the assignment of a general lower bound as $\Delta F_{\textrm{effective}}$ for the average work or released heat associated with the erasure mechanisms. This outcome of the study bears a powerful significance. This is because it allows us to predict a limiting value for $\langle W\rangle$ or $\langle Q\rangle$ for the cases even when the classical Landauer bound does not hold. This departure from the Landauer limit can happen due to the non-idealized nature of the memory states, as in our present cases. We emphasize that it is possible to propose appropriate designs of the systems that can improve the operations of the erasure mechanisms in terms of the thermodynamic costs when the initial memory states have an asymmetric equilibrium or nonequilibrium origin. However, we retrieve the Landauer bound under the conventional setup with initially symmetric and equal distribution of memory states. This comparative study gives us the scope of proper quantitative understanding of improving the suggested erasure protocols from the perspective of energetics. Also, our analyses unveil another very vital fact that the success of the erasure process can be enhanced through the tuning of the asymmetry factor governing the dynamics. Therefore, it can be considered that the current work defines a systematic path and develops a detailed comprehension for the improvement of the efficiency of the fundamental process of erasure. We believe that the knowledge acquired from the present study would guide the design of more efficient erasure mechanisms. Here, we also mention that the analyses and the results of the current work suggest the possible ways of improvements for the erasure process, specifically. We can expect that the construction of improved erasure mechanisms may upgrade the overall performance of the digital devices. This is because erasure is an essential step for irreversible computation, measurement (resetting to the initial memory state step), and communication, the primary functions in digital devices [6]. However, there are several other steps involved in addition to the erasure mechanism in these three basic operations (computation, measurement and communication) occurring inside the digital apparatus. Therefore, we indicate that our present study cannot conclusively suggest the complete upgradation of the performance of the digital machines, as there are many other factors that influence the functions of such devices. However, it definitely offers a useful scheme, employing asymmetric memory states, for the improvement of the functions of one common essential step, the erasure mechanism, which is involved in all three fundamental tasks in digital devices. We also point out that the focus on this particular step is very important, as this accounts for the heat release in the environment during the devices’ actions. However, to obtain an integrated understanding of the overall level of execution of the devices, precise scrutiny of the performance of each individual step in the information processing systems, considering asymmetric memory states, will be required. These explorations can be considered as extensive points of research in future to aim at designing improved technological devices.